Kira wanted to stock up on drinks for an upcoming party. First she spent $53 on 14 cases of juice and 11 cases of soda, which is all the store had in stock. A few days later, she returned to the store and purchased an additional 14 cases of juice and 8 cases of soda, spending a total of $50. What is the price of each drink?
Found 3 solutions by josgarithmetic, ikleyn, MathTherapy:
Answer by josgarithmetic(39630)

If first and second purchases gave a total of 50 dollars, then the second purchase cost was 50-53=-3 which makes no sense.

Say more specifically, that the FIRST purchase was $53 and the SECOND purchase was $50. Now, this total IS $103; and their difference is $3.

Answer by ikleyn(52905)

.
Kira wanted to stock up on drinks for an upcoming party.
First she spent $53 on 14 cases of juice and 11 cases of soda, which is all the store had in stock.
A few days later, she returned to the store and purchased an additional 14 cases of juice and 8 cases of soda, spending a total of $50.
What is the price of each drink?
~~~~~~~~~~~~~~~~

This problem is for mental solution.

1.  After reading the condition, you can easily see that the difference $53 - $50 = $3 is exactly the cost of 11-8 = 3 cases of soda,

    so that one case of soda costs 1 dollars.


2.  After finding this, you conclude that 14 cases of juice cost  53 - 11*1 = 42 dollars.  Hence, each case of the juice costs  42/14 = 3 dollars.

Solved.

--------------
To see other similar solved problems,  look into the lesson
    - Solving mentally word problems on two equations in two unknowns
in this site.

--------------

Ignore writing by @josgarithmetic,  because it is nonsense.



Answer by MathTherapy(10557)


Kira wanted to stock up on drinks for an upcoming party. First she spent $53 on 14 cases of juice and 11 cases of soda, which is all the store had in stock. A few days later, she returned to the store and purchased an additional 14 cases of juice and 8 cases of soda, spending a total of $50. What is the price of each drink?
Correct answer: